OpenStudy (anonymous):

Zach bought a cheeseburger and bottle of water for $9. Jasmine got an order of nachos and a bottle of water for $8. Mr. Johnson purchased five bottles of water for his family for $15. Write equations for each purchase. Solve your equations to find the cost of a bottle of water, order of nachos and a cheeseburger

Directrix (directrix):

Let C represent one cheeseburger; W, one bottle of wate and, N, one order of nachos. (1) C + W = 9 (2) N + W = 8 (3) 5W = 15 ---> W = 3 Back-substituting into (2), N + 3 = 8, and N = 5 Back-substituting into (3), C + 3 = 9, and C = 6.

OpenStudy (anonymous):

start with 5 bottles of waters $15